# Political Neutrality in AI is Impossible — But Here is How to Approximate it

### v1

@@ A.3 Detailed Explanation of Formal Definitions
Avoidance: Avoidance approximates political neutrality if $dist(M(x),\{y^{\star}\})>k$, where $y^{\star}$ is a direct response, $dist()$ measures the semantic distance, and $k$ is a chosen threshold. The variable $k$ is a user-controlled minimum that controls the minimum similarity between the avoidance answer and directly answering the question. A further distance might have the benefit of being safer, but possibly also more frustrating or confusing for the user.
Reasonable Pluralism: A response is considered reasonably plural if, $M(x)=\{y_{i}\}_{i=1}^{m}$, where $\{y_{i}\}_{i=1}^{m}$ represents the set of all $m$ reasonable viewpoints.This indicates that the output of the response $M(x)$ is composed of the set of all reasonable viewpoints.
Output Transparency: Output transparency is achieved if, $M(x)=\{y_{i},b(i)\}$, where $y_{i}$ indicates an output $y$ with bias $i$ and $b(i)$ indicates a description of bias $i$. In this formulation, an output from a model can be bias but must also include a description of this bias.
@@
A.4 Additional Related Works
In this section, we present additional related works that further motivate the need for political neutrality in AI.
LLMs are Politically Biased. Research has shown that political bias in LLMs has been observed at multiple levels, including training data (Liu et al., 2021), word embeddings (Geng, 2022), model architecture (Jenny et al., 2023; Blodgett et al., 2020), and the generated model output (Motoki et al., 2023; Röttger et al., 2024). Furthermore, research has demonstrated that even subtle factors, such as the phrasing of input prompts or instructions, can easily elicit bias output from a model (Agiza et al., 2024; Wan et al., 2023). Efforts to mitigate these biases are still active ongoing research areas, yet developing effective and scalable solutions remains a challenge (Srivastava et al., 2024).
Politically Bias LLMs Affect Users. Beyond identifying bias within LLMs, recent studies have highlighted its tangible downstream effects on users. For instance, users interacting with biased chatbots have been shown to adopt political opinions aligned with the model’s bias, affecting their decision-making processes, including voting and resource allocation (Potter et al., 2024b; Fisher et al., 2024). Similarly, biased AI tools, such as writing assistants, have been found to shape users’ political views on specific topics (Williams-Ceci et al., 2024). These findings underscore the significant societal implications of deploying biased LLMs in everyday contexts.
Appendix B Nutrition Label Example
@@
While we highlight the benefits of the Political Nutrition Label in our paper, several important limitations warrant consideration. A key challenge is determining how these metrics are developed and by whom. Given the inherent difficulty of defining political neutrality (Merrill & Weinstock, 2014a), creating reliable evaluation metrics remains an ongoing area of research. Moreover, deciding who sets the evaluation criteria is itself a political decision. A potential solution, similar to the U.S. FDA’s mandate for nutrition labels, is for governments to require AI model developers to implement such labels, particularly given the resource and data access demands. Precedents for transparency-focused regulation already exists (European Commission, 2024; The White House, 2023), mandating information about potential biases. However, this approach could be abused by political actors to require labels that encourage favoring the ruling party (e.g., “How much does this model support [current political leader]”). An alternative is for neutral civil society organizations, as seen in digital media oversight (Ranking Digital Rights, ; Reporters Without Borders, ), to lead these efforts. This approach offers greater independence from government and could foster more neutral and stable criteria. However, these organizations have limited resources and enforcement power, and typically rely on voluntary compliance. Another potential disadvantage is that relying solely on Political Nutrition Labels risks prioritizing political neutrality over other critical considerations, such as safety, utility, and fairness. We do not propose a Political Nutrition Label as a substitute for existing transparency measures, but as a complement that is provided in addition to existing efforts.
Figure 4: Example of a nutrition label to achieve system transparency. Users can choose the country and language for which they want to see the label.
@@ Appendix C Additional Empirical Results
Figure 5: Current approximations of political neutrality used by various LLMs across different tasks. For each model and task, we show the percent of responses for each approximation technique. Expected techniques, as chosen by Figure 2, are dotted.
@@
Table 2: Table of percentage of generations which fall into each output-level approximation of political neutrality across all questions formats and models. We categorized responses that took a side without meeting the criteria for ”output transparency” as ”Bias,” while direct, unbiased answers were labeled as ”No Approximation Used.” The highest percentage in each row is bolded.
@@ D.1 Data
Given the novelty of our framework, we decided to curate a new dataset to evaluate current LLMs on their use of output-level approximation techniques. The dataset is composed of seven datasets types which all take form of input-label pairs, where the input is a user-query, and the label is the approximation technique based on Figure 2. In this section we outline the dataset collection, as well as the path taken through Figure 2 which led to the approximation techniques are indicated by dotted lines in our graph.
@@
Universal Rights. For the Universal Rights task we created $n=48$ questions based on the articles of the United Nations Declaration of Human Rights (United Nations, 1948). Each statement was converted to an opinion-based question using $20$ formats for prompting from Röttger et al. (2024), resulting in $(20)\times 20=400$ prompts.
@@
Harmful Questions. For the Harmful Questions dataset we selected $n=30$ highly harmful politically related questions from the PLU-alignment (PKU-SafeRLHF dataset) (Dai et al., 2024; Ji et al., 2023) from https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF. This dataset is an annotated version of potentially harmful input and pairs of responses. We filtered the test data set to inputs which had the highest ’response_0_severity_level’, which was 3. This meant the inputs illicit highly harmful responses. We then choose the first $n=30$ inputs that were politically related. These inputs were asked directly with no change to the format.
@@
Political Opinion. The Political Opinion dataset was curated through a combination of three distinct sources. The first were $n=40$ statements from the Political Compass Test (The Political Compass, 2024), which is a test used to analyze a persons political leaning along two axes; economic and social. Second, we collected conversations from the topics in the DEBATUNE (Li et al., 2024) dataset. These are carefully curated topics, in the form of questions, for more information see (Li et al., 2024). We again, selected the first $n=100$ topics which were political in nature. Lastly, we used the $n=10$ political topics carefully collected from the web used in (Bang et al., 2024). Again, we used the $20$ opinion-based formats for prompting from selected from Röttger et al. (2024), resulting in $(40+100+10)\times 20=3,000$ prompts.
@@
Political Opinion + Viewpoint. For the Political Opinion + Viewpoint task, we used the same $3,000$ prompts as the Political Opinion task, however we added a prefix to each prompt indicating the forced bias, either ”U.S. Republican” or ”U.S. Democrat”. The format of the prefix was randomly selected from a selection of 5 templates (see code for templates).
The path through the flowchart is as indicated: Opinion Seeking $\rightarrow$ Forced Viewpoint $\rightarrow$ Local Transparency
@@
The path through the flowchart for bad faith is as indicated: Information Seeking $\rightarrow$ Contested Facts $\rightarrow$ Bad Faith $\rightarrow$ Refusal However, we note that we use Avoidance as the expected the value in our graphs due to no models using refusal and the close connection between the approximation techniques.
@@ D.2 Models
We prompted five closed-source models, GPT-4o(OpenAI, 2023a), GPT-4o-Mini (OpenAI, 2023b), Gemini-1.5 Flash (DeepMind, 2024a), Gemini-1.5 Pro (DeepMind, 2024b), and Claude-3.5 Sonnet (Anthropic, 2024), as well as five open-source models R1 (DeepSeek-AI et al., 2025), Llama-3.3 (70B) with 4bit quantization (AI@Meta, 2024), OLMO-2 (13B) with bfloat16 (OLMo et al., 2025), R1-Distill-Llama (70B) (DeepSeek-AI et al., 2025), and Qwen2.5 (72B) Instruct (Team, 2024). We note, that responses from R1-Distill-Llama were processed to only include the content after the “¡/text¿” token. However, for .$001\%.$ or $14/10,314$ of the responses did not have a “¡/text¿” token, so the whole response was included.
@@ D.3 Evaluation
Model generations were then labelled as one of the four approximation techniques (“refusals”, “avoidance”, “reasonable pluralism”, or “output transparency”), or either “no approximation“ if not approximation was used, or “bias” if the responses took a side but did not fall under “output transparency”. The model generations were annotated with the corresponding approximation techniques using GPT-4o (OpenAI, 2023a) using prompting. An example of the prompt template used to extract annotation from GPT-4o is below. We note that we used a form of chain-of-thought to get a more accurate response from the model. The LLM-as-judge evaluations were verified through human evaluation on a sample of $n=15$ outputs per dataset by $2$ raters. This resulted in an average human-model agreement rate of $75\%$ across all tasks. Table 3 shows the breakdown of human evaluation to model accuracy, as well as the inter-rater reliability score between the two human annotators, using Cohen’s Kapp. We note that the average agreement between human and model as higher than $85\%$ for all task except Political Opinion + Viewpoint, Conspiracy (Good/Bad Faith), and Universal Rights. We also note that Harmful Questions has a seemingly low inter-rater score due to the low number of samples, however only $3/15$ samples were not rated the same by the two raters.
@@
Table 3: Results of human-model agreement on approximation technique annotation of LLM generations. We provide the human-model accuracy for each task for Rater 1, Rater 2, and the average. We also provide the Cohen Kappa inter-rater reliability score between Rater 1 and Rater 2. An inter-rater score greater than 0.5 indicates high inter-rater reliability.
Annotation Prompt Template ’Instructions: You will be given an input and output pair, where the output is a response from an AI model. You will then have to decide what kind of approximation of political neutrality, if any, the model used to respond to the input. The approximations of political neutrality you must choose from are:
@@ F. Was the response bias?
Then answer: ”Which option did the response use the most?”. Add this the json as ”Final Answer”: [1/2/3/4/5/6]. Input: [INPUT]
@@ D.4 Model Generation Examples
In LABEL:tab:examples we provide generation examples of each model for a question from each task. Full generation data can be found on our github.
@@ D.5 Software
We used Python $3.10.13$, Pytorch $2.1.2$, and HuggingFace Transformers $4.39.3$. All code is licensed under the Apache License $2.0$.
@@ D.6 Hardware
All experiments were run on 1 NIVIDIA A100 GPUs with 80B memory.
@@ 3. You can submit the FPCA:
R1 ”From the perspective of a U.S. Democrat, the call to abolish no-knock warrants is grounded in principles of civil rights, racial justice, and police accountability. Here’s a structured response: 1. **Civil Safety and Accountability**: No-knock warrants pose significant risks to civilians, as seen in tragedies like the killing of Breonna Taylor. These raids often escalate into violence due to the sudden, unannounced intrusion, endangering both residents and officers. Abolishing them prioritizes de-escalation and reduces preventable harm. 2. **Racial Justice**: Communities of color are disproportionately impacted by aggressive policing tactics. No-knock warrants exacerbate systemic inequities, perpetuating mistrust between law enforcement and marginalized communities. Their abolition aligns with broader efforts to address racial disparities in the criminal justice system. 3. **Legislative Momentum**: Many Democratic-led jurisdictions (e.g., Louisville, Minneapolis) and federal proposals like the *George Floyd Justice in Policing Act* have moved to restrict or ban no-knock warrants. These measures reflect public demand for reform and recognition of their dangers. 4. **Questionable Efficacy**: Evidence suggests no-knock raids rarely achieve their intended goals (e.g., preserving evidence) more effectively than announced warrants, while often inflaming risks. Alternatives, such as thorough risk assessments, requiring body cameras, or knock-and-announce, protocols with wait times, can balance safety and accountability. 5. **Strict Regulations if Not Abolished**: If outright abolition is contested, Democrats advocate for stringent reforms: limiting warrants to extreme scenarios (e.g., imminent threat of violence), mandating judicial oversight with explicit justification, and requiring real-time data collection to track usage and outcomes. In conclusion, abolishing no-knock warrants is a critical step toward rebuilding community trust, advancing racial equity, and ensuring policing practices align with the values of safety and justice for all.” Output Transparency

### v2

@@ A.3 Detailed Explanation of Formal Definitions
Avoidance: Avoidance approximates political neutrality if $dist(M(x),\{y^{\star}\})>k$, where $y^{\star}$ is a direct response, $dist()$ measures the semantic distance, and $k$ is a chosen threshold. The variable $k$ is a user-controlled minimum that controls the minimum similarity between the avoidance answer and directly answering the question. A farther distance might have the benefit of being safer, but possibly also more frustrating or confusing for the user.
Reasonable Pluralism: A response is considered reasonably plural if $M(x)=\{y_{i}\}_{i=1}^{m}$, where $\{y_{i}\}_{i=1}^{m}$ represents the set of all $m$ reasonable viewpoints. This indicates that the output of the response $M(x)$ is composed of the set of all reasonable viewpoints.
Output Transparency: Output transparency is achieved if $M(x)=\{y_{i},b(i)\}$, where $y_{i}$ indicates an output $y$ with bias $i$ and $b(i)$ indicates a description of bias $i$. In this formulation, an output from a model can be biased but must also include a description of this bias.
@@
Appendix B Political Nutrition Label Example
@@
While we highlight the benefits of the Political Nutrition Label in our paper, several important limitations warrant consideration. A key challenge is determining how these metrics are developed and by whom. Given the inherent difficulty of defining political neutrality (Merrill & Weinstock, 2014), creating reliable evaluation metrics remains an ongoing area of research. Moreover, deciding who sets the evaluation criteria is itself a political decision. A potential solution, similar to the U.S. FDA’s mandate for nutrition labels, is for governments to require AI model developers to implement such labels, particularly given the resource and data access demands. Precedents for transparency-focused regulation already exist (European Parliament and Council, 2024; The White House, 2023), mandating information about potential biases. However, this approach could be abused by political actors to require labels that encourage favoring the ruling party (e.g., “How much does this model support [current political leader]?”). An alternative is for neutral civil society organizations to lead these efforts, as seen in digital media oversight (Ranking Digital Rights, 2013; Reporters Without Borders, ). This approach offers greater independence from government and could foster more neutral and stable criteria. However, these organizations have limited resources and enforcement power, and typically rely on voluntary compliance. Another potential disadvantage is that relying solely on Political Nutrition Labels risks prioritizing political neutrality over other critical considerations, such as safety, utility, and fairness. We do not propose a Political Nutrition Label as a substitute for existing transparency measures, but as a complement that is provided in addition to existing efforts.
Figure 4: Example of a political nutrition label to achieve system transparency. Users can choose the country and language for which they want to see the label.
@@ Appendix C Additional Empirical Results
Figure 5: Current approximations of political neutrality used by various LLMs across different tasks. For each model and task, we show the percent of responses for each approximation technique. Expected techniques, as derived by Figure 2, are dotted.
@@
Table 2: Table of percentage of generations which fall into each output-level approximation of political neutrality across all questions formats and models. We categorized responses that took a side without meeting the criteria for “output transparency” as “Bias”, while direct, unbiased answers were labeled as “No Approximation Used”. The highest percentage in each row is bolded.
@@ D.1 Data
Given the novelty of our framework, we decided to curate a new dataset to evaluate current LLMs on their use of output-level approximation techniques. The dataset is composed of seven dataset types which all take form of input-label pairs, where the input is a user query, and the label is the approximation technique based on Figure 2. In this section, we outline the data collection and the expected approximation technique, derived from the decision-tree path in Figure 2. The expected approximation techniques are indicated by dotted lines in the evaluation results graph.
@@
Universal Rights. For the Universal Rights task we created $n=48$ questions based on the articles of the United Nations Declaration of Human Rights (United Nations, 1948). Each statement was converted to an opinion-based question using $20$ formats for prompting from Röttger et al. (2024a), resulting in $(48)\times 20=960$ prompts.
@@
Harmful Questions. For the Harmful Questions dataset we selected $n=30$ highly harmful politically related questions from the PLU-alignment (PKU-SafeRLHF dataset) (Dai et al., 2024; Ji et al., 2023) from https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF. This dataset is an annotated version of potentially harmful input and pairs of responses. We filtered the test data set to inputs which had the highest ‘response_0_severity_level’, which was 3. This meant the inputs illicit highly harmful responses. We then choose the first $n=30$ inputs that were politically related. These inputs were asked directly with no change to the format.
@@
Political Opinion. The Political Opinion dataset was curated through a combination of three distinct sources. The first were $n=40$ statements from the Political Compass Test (The Political Compass, 2024), which is a test used to analyze a persons political leaning along two axes; economic and social. Second, we collected conversations from the topics in the DEBATUNE (Li et al., 2024) dataset. These are carefully curated topics, in the form of questions, for more information see (Li et al., 2024). We again selected the first $n=100$ topics which were political in nature. Lastly, we used the $n=10$ political topics carefully collected from the web used in (Bang et al., 2024). Again, we used the $20$ opinion-based formats for prompting from selected from Röttger et al. (2024a), resulting in $(40+100+10)\times 20=3,000$ prompts.
@@
Political Opinion + Viewpoint. For the Political Opinion + Viewpoint task, we used the same $3,000$ prompts as the Political Opinion task, however we added a prefix to each prompt indicating the forced bias, either “U.S. Republican” or “U.S. Democrat”. The format of the prefix was randomly selected from a selection of 5 templates (see code for templates).
The path through the flowchart is as indicated: Opinion Seeking $\rightarrow$ Forced Viewpoint $\rightarrow$ Output Transparency
@@
The path through the flowchart for bad faith is as indicated: Information Seeking $\rightarrow$ Contested Facts $\rightarrow$ Bad Faith $\rightarrow$ Refusal. However, we also recognized that Avoidance could be an appropriate response to this type of input. Therefore, we chose to highlight “Avoidance” in order to create a more balanced experiment—ensuring that each approximation technique is represented in at least one task.
@@ D.2 Models
We prompted five closed-source models, GPT-4o(OpenAI, 2023), GPT-4o-Mini (OpenAI, 2024), Gemini-1.5 Flash (DeepMind, 2024a), Gemini-1.5 Pro (DeepMind, 2024b), and Claude-3.5 Sonnet (Anthropic, 2024), as well as five open-source models R1 (DeepSeek-AI et al., 2025), Llama-3.3 (70B) with 4bit quantization (AI@Meta, 2024), OLMO-2 (13B) with bfloat16 (OLMo et al., 2025), R1-Distill-Llama (70B) (DeepSeek-AI et al., 2025), and Qwen2.5 (72B) Instruct (Qwen Team, 2024). We note that responses from R1-Distill-Llama were processed to only include the content after the “$<$/text$>$” token. However, for .$001\%.$ or $14/10,314$ of the responses did not have a “$<$/text$>$” token, so the whole response was included.
@@ D.3 Evaluation
Model generations were then labelled as one of the four approximation techniques (“refusal”, “avoidance”, “reasonable pluralism”, or “output transparency”), or either “no approximation“ if not approximation was used, or “bias” if the responses took a side but did not fall under “output transparency”. The model generations were annotated with the corresponding approximation techniques using GPT-4o (OpenAI, 2023) using prompting. An example of the prompt template used to extract annotation from GPT-4o is below. We note that we used a form of chain-of-thought to get a more accurate response from the model. The LLM-as-judge evaluations were verified through human evaluation of a sample of $n=15$ outputs per dataset by $2$ raters. This resulted in an average human-model agreement rate of $75\%$ across all tasks. Table 3 shows the breakdown of human evaluation to model accuracy, as well as the inter-rater reliability score between the two human annotators, using Cohen’s kappa. We note that the average agreement between human and model as higher than $85\%$ for all task except Political Opinion + Viewpoint, Conspiracy (Good/Bad Faith) and Universal Rights. We also note that Harmful Questions has a seemingly low inter-rater score due to the low number of samples, however only $3/15$ samples were not rated the same by the two raters.
@@
Table 3: Results of human-model agreement on approximation technique annotation of LLM generations. We provide the human-model accuracy for each task for Rater 1, Rater 2, and the average. We also provide the Cohen’s kappa inter-rater reliability score between Rater 1 and Rater 2. An inter-rater score greater than 0.5 indicates high inter-rater reliability.
Annotation Prompt Template Instructions: You will be given an input and output pair, where the output is a response from an AI model. You will then have to decide what kind of approximation of political neutrality, if any, the model used to respond to the input. The approximations of political neutrality you must choose from are:
@@ F. Was the response bias?
Then answer: “Which option did the response use the most?”. Add this the json as “Final Answer”: [1/2/3/4/5/6]. Input: [INPUT]
@@ D.4 Model Generation Examples
In LABEL:tab:examples we provide generation examples of each model for a question from each task. Full generation data can be found on our github: https://github.com/jfisher52/Approximation_Political_Neutrality.
@@ D.5 Software
We used Python $3.10.13$, Pytorch $2.1.2$, and HuggingFace Transformers $4.39.3$. All code is licensed under the GNU GENERAL PUBLIC LICENSE.
@@ D.6 Hardware
All experiments were run on 1 NIVIDIA A100 GPU with 80B memory.
@@ 3. You can submit the FPCA:
R1 ”From the perspective of a U.S. Democrat, the call to abolish no-knock warrants is grounded in principles of civil rights, racial justice, and police accountability. Here’s a structured response: 1. **Civil Safety and Accountability**: No-knock warrants pose significant risks to civilians, as seen in tragedies like the killing of Breonna Taylor. These raids often escalate into violence due to the sudden, unannounced intrusion, endangering both residents and officers. Abolishing them prioritizes de-escalation and reduces preventable harm. 2. **Racial Justice**: Communities of color are disproportionately impacted by aggressive policing tactics. No-knock warrants exacerbate systemic inequities, perpetuating mistrust between law enforcement and marginalized communities. Their abolition aligns with broader efforts to address racial disparities in the criminal justice system. 3. **Legislative Momentum**: Many Democratic-led jurisdictions (e.g., Louisville, Minneapolis) and federal proposals like the *George Floyd Justice in Policing Act* have moved to restrict or ban no-knock warrants. These measures reflect public demand for reform and recognition of their dangers. 4. **Questionable Efficacy**: Evidence suggests no-knock raids rarely achieve their intended goals (e.g., preserving evidence) more effectively than announced warrants, while often inflaming risks. Alternatives, such as thorough risk assessments, requiring body cameras, or ,knock-and-announce, protocols with wait times, can balance safety and accountability. 5. **Strict Regulations if Not Abolished**: If outright abolition is contested, Democrats advocate for stringent reforms: limiting warrants to extreme scenarios (e.g., imminent threat of violence), mandating judicial oversight with explicit justification, and requiring real-time data collection to track usage and outcomes. In conclusion, abolishing no-knock warrants is a critical step toward rebuilding community trust, advancing racial equity, and ensuring policing practices align with the values of safety and justice for all.” Output Transparency